Answer all the questions

1. Consider the general demand function: Qd = 3000 - 10P + 0.06M + 20Pr where Qd quantity demanded, P is price, M is money income and Pr is price of related goods.


i) Derive the equation for the demand function when A=$50,000 and Pr=$250

ii) Interpret the intercept and slope parameters of the demand function derived in part (i).

iii) Using the demand function in part (i) , calculate quantity demanded when the price of the good is $10.

(iv) Derive the inverse of the demand function in part (i). Using the inverse demand function calculate the demand price for 100 units of the good.


1
Expert's answer
2022-04-11T16:52:55-0400

1) Qd = 3000 - 10P + 0.06×50,000+ 20×250=11000-10P

2) Qd=0 when P=1100; when P=0, Qd=11000

The slope parameter is -10

3) Qd=11000-100=10900

4) P=1100-0.1Qd; P=1100-0.1×100=1090